OpenStudy (anonymous):

how wold I graph this? 1 -3/2 x +4.

OpenStudy (anonymous):

Slope of y= -3/2 x +4.2 is -3/2 x=0,y=4 y-intercept is 4

sam (.sam.):

y-intersect is your 'c' in y=mx+c

sam (.sam.):

so y-intersect =4.2 slope is correct when x=0, y=4.2